Download
operation tcr n.
Skip this Video
Loading SlideShow in 5 Seconds..
Operation TCR PowerPoint Presentation
Download Presentation
Operation TCR

Operation TCR

138 Vues Download Presentation
Télécharger la présentation

Operation TCR

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. Operation TCR 2nd of July to 10th of September 2001 More than 20 major events Events grouped by equipment group and zone affected Niels.Schionnemann@Cern.ch

  2. Electrical events • 19/08 18:42: Electrical disturbance of 400 kV EDF due to storm, main PS magnet tripped. No electrical equipment switched off. • 19:15 PS running again • 19/08 21:08: Compensator 1 BE tripped due to storm. SPS and PS stopped. • 21:21 PS running again • 21:40 Meyrin filters fall out. PS still ok • 22:14 Compensator in BE ok, SPS running • 22:29 Meyrin filters ok • 20/08 00:34: Compensator 1 BE tripped due to storm. PS and SPS stopped • 1:30 PS and SPS running again Niels Schionneman ST/MO/TCR

  3. Electrical events • 20/8 06:48 Short circuit in transformer EHT3 BE • Transfo was idling but the short circuit caused an unbalance in the network and made the compensators trip. • Demi water off Prevessin and Meyrin • PS and SPS stopped • Electrical problem solved at 7:20 • Cooling PS ok at 9:20 • Cooling SPS ok at 10:20 Niels Schionneman ST/MO/TCR

  4. Electrical events • 29/8 18:47 Explosion in 18kV breaker sub-station 7 • 7 18kV breakers in the PS loop trip • The 66kV supply for the transfo SEH1 tripped. Meyrin off tension • Auto transfer switched to EOS after 5 sec • The 18kV trips due to malfunction of protection equipment/selectivity problem • PS and SPS stopped for 24h Niels Schionneman ST/MO/TCR

  5. CV events concerning PS • 3/7 14:20 Water leak alarm in PS • Demanded access but found no leak. Drains were blocked and condense water from air condition caused the water level to raise • 15:33 PS running again • 4/7 21:43 Cooling problem in 361 • MCR can not switch on main magnet due to high temperature in 361 • Ventilator in cooling tower not running on max speed • 23:43 PS running again Niels Schionneman ST/MO/TCR

  6. CV events concerning PS • 15/7 9:01 PS demi water stopped • A UPS supplying the PLCs for the demi water stations had tripped due to water infiltration • UPS bypassed and PLCs and water stations started • 11:25 PS running again • 19/8 2:43 and 7:19 Demi water station 359 stopped. PS stopped • Station stopped twice due to heavy storms. Started station manually • 3:05 and 7:39 Cooling ok Niels Schionneman ST/MO/TCR

  7. CV events concerning PS • 20/8 19:00 Demi water for TT2 stopped • No beam for SPS • No alarms received in TCR • Magnet in TT2 supplied from TOF line. To be changed • 20:20 Cooling station in 378 started again • 23/8 20:15 Demi water temperature too high in East area • Pumping station running on max • Primary water temperature at 25 degrees due to very warm weather. • Informing MCR. Situation unchanged all night Niels Schionneman ST/MO/TCR

  8. CV events concerning PS • 25/8 15:10 Pump in cooling circuit for ISOLDE exploded • ISOLDE stopped • No fluid alarms in TCR, only electrical alarm • Isolated pump and started spare pump • 15:40 Cooling ok Niels Schionneman ST/MO/TCR

  9. CV events concerning SPS • 6/7 17:45 Temperature interlock on main magnet in BA3. SPS stopped • Supply valve closed on primary side of heat exchanger • 19:38 SPS running again • 9/7 11:45 Temperature interlock on septum magnet in BA6. SPS stopped • Pump stopped by faulty pressure sensor • 12:30 SPS running again • 27/7 10:25 BA6 cooling off. SPS stopped • CV working on PLC for BA6 cooling. A faulty Siemens module caused the station to stop • 10:55 SPS running again Niels Schionneman ST/MO/TCR

  10. CV events concerning SPS • 2/8 14:27 Demi water BA81 stopped due to water leak. No beam for North area • Access to machine given but leak is not found • Manages to start pumps and refill continuously • 15:39 North area beam back on • 5/8 Compensator tripped due to missing cooling in BA3. SPS Stopped • No alarms in TCR • Cooling stopped due to problem with the PLC • 19:40 SPS running again • 10/8 20:40 Same problem as above • 21:30 SPS running again Niels Schionneman ST/MO/TCR

  11. CV events concerning SPS • 27/8 4:20 Cooling for septum magnet BA2 stopped. SPS stopped • No alarms in the TCR. Pump restarted from TCR • 4:22 SPS running again Niels Schionneman ST/MO/TCR

  12. CV events concerning SPS • 2/9 4:31 + 3/9 4:56 Demi water BA3 stopped. No alarms in TCR • Problem seen when compensator in BB3 tripped • SPS stopped for 1h each time • Cooling stopped due to error in PLC • 3/9 23:59 Demi water for CAV2+RF1 BB3 stopped • SPS stopped for 3.5h • Cooling stopped due to electrical problem on pump for CAV2 circuit. Pumps started in forced mode. Niels Schionneman ST/MO/TCR

  13. Various events • 14/7-15/7 RTserver crashes several times • Different zones at Meyrin without supervision for various periods. • Fire brigade instructed to do rounds and keep a close eye on the synoptic. • 16/7 Bugs found and fixed by IN. Situation back to normal. • 27/8 3:55 Vacuum sector valve closed TS1 and TS2. SPS stopped. • 4:00 Valve in TS1 opened. OK for SPS • 4:10 Valve in TS2 closes • 4:12 Valve in TS2 opened. OK for SPS Niels Schionneman ST/MO/TCR